OpenStudy (anonymous):

What is the slope of a line parallel to 4x + y = -1? A) m = -4 B) m = 2 C) m = -2 D) m = -5 E) m = -3

OpenStudy (anonymous):

What is the slope of a line parallel to 10x - 5y = 8? What is the slope of a line parallel to 4x + 2y = 5? What is the slope of a line perpendicular to x - 5y = -10? What is the slope of a line perpendicular to x - 3y = 9? The answers to these questions are the a,b,c,d,e from above.

OpenStudy (anonymous):

m = -4

OpenStudy (anonymous):

slope of a line parallel to 10x - 5y = 8 10x_8=5y y= 2x-8/5 Slope m= 2

OpenStudy (anonymous):

slope of a line parallel to 4x + 2y = 5 slope m= -2

OpenStudy (anonymous):

slope of a line perpendicular to x - 5y = -10 i.e. 5y = x+10 i.e. y=1/5x+10 Slope of the required line m =-5

OpenStudy (anonymous):

slope of a line perpendicular to x - 3y = 9 i.r. x-9 =3y i.e. y= 1/3x-3 slope of the required line m= -3

OpenStudy (anonymous):

slope of a line perpendicular to x - 3y = 9 i.r. x-9 =3y i.e. y= 1/3x-3 slope of the required line m= -3

OpenStudy (anonymous):

Thanks @dpasingh!